In its first month of operations, Literacy for the Illiterate opened a new bookstore and bought merchandise in the following order: (1) 200 units at $7 on January 1, (2) 500 units at $8 on January 8, and (3) 800 units at $9 on January 29.

Assume 975 units are on hand at the end of the month, calculate the cost of goods available for sale, ending inventory, and cost of goods sold under the weighted average cost flow assumptions. Assume a periodic inventory system is used.

(Round "Cost per Unit" to 2 decimal places.)

Weighted Average Cost Cost of Goods Available for Sale Ending Inventory Cost of Goods Sold
Answer #1
Units Unit cost Total
January 1 200 7 1400
January 8 500 8 4000
January 29 800 9 7200
Total 1500 12600
Average cost per unit 8.40 =12600/1500
Cost of goods available for sale 12600
Ending inventory 8190 =975*8.40
Cost of goods sold 4410 =(1500-975)*8.40